HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

IMPROVED THEORY OF THE EFFECTIVE DIPOLE MOMENT AND ABSOLUTE LINE STRENGTHS OF THE XY_{2} ASYMMETRIC TOP MOLECULE IN THE X^{2}B_{1} DOUBLET ELECTRONIC STATE

\underline{O. N. ULENIKOV} , \underline{E. S. BEKHTEREVA} , \underline{O. V. GROMOVA} , \underline{A. N. KAKAULIN}, National Research Tomsk Polytechnic University, 30, av.

Lenina, 634050 Tomsk, Russia; \underline{C. SYDOW} and \underline{S. BAUERECKER} , Institut für Physikalische und Theoretische Chemie, Technische Universität Braunschweig, D -

38106, Braunschweig, Germany

A new effective dipole moment model for the XY_{2} (C_{2v}−symmetry) molecule in a doublet electronic state is derived that includes (as special cases) all currently known models of effective dipole moment for such type molecules, and allows, at the same time, to take into account an influence of spin–rotation interactions on the effective dipole moment operator which were not taken into account in the before studies.

The study was financially supported by the Russian Science Foundation, project 22–22–00171.

p-number: p014  Submitted on Sat, 27 May 2023 14:36:57 +0200